OpenStudy (anonymous):

If n+2/5 is an integer, what is the remainder when n is divided by 5?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

n could equal any number

OpenStudy (anonymous):

3/5 + 2/5 = 5/5 = 1 (integer) remainder is 3 8/5 + 2/5 = 10/5 = 2 (integer) remainder is 3 whichever n that works for the equation to be an integer has remainder 3 when divided by 5.

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

n=1

OpenStudy (anonymous):

1+2/5=3/5

OpenStudy (anonymous):

where is the remainder of 3?

OpenStudy (anonymous):

3/5 is no integer

OpenStudy (anonymous):

ahhh, derp

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

but don't they mean 3/5

OpenStudy (anonymous):

3 divided by 5

OpenStudy (anonymous):

n=3

OpenStudy (anonymous):

3 divided by 5

OpenStudy (anonymous):

3/5+2/5=5/5

OpenStudy (anonymous):

3/5 = 0 and the remainder is 3

OpenStudy (anonymous):

1

OpenStudy (anonymous):

3/5=.6

OpenStudy (anonymous):

8+2/5=10/5=2

OpenStudy (anonymous):

yes, but if they asked for remainder you don't get to the decimal part

OpenStudy (anonymous):

so what you are trying to say is?

OpenStudy (anonymous):

13/5 + 2/5 = 15/5 = 3 13/5 = 2 r 3

OpenStudy (anonymous):

8/5 gives remainder 3

OpenStudy (anonymous):

3/5 gives remainder 3?

OpenStudy (anonymous):

Yes, the reaminder is going to be 3.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so a fraction can have a remainder?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

3/5 has a remainder of 3?

OpenStudy (anonymous):

A fraction doesn't have remainder, but a number does, they didn't ask for a remainder for n/5, only they asked a remainder for n. 3/5, the n = 3. The remainder of 3 when divided by the number 5 is 3. And so is the remainder for all values of n allowed if the equation must be an integer. n = 3, 8, 13, 18,..

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Yw